OpenStudy (anonymous):

An office used 200 stamps. Some were .35 cents and some were .20cents. If the total was 62.95, how many did the office use?

OpenStudy (anonymous):

set up a system of equations (pretty sure this is them) x+y=200 .35x+.20y=62.95 then solve via the elimination or substatution methode

OpenStudy (anonymous):

Solve the following for x and y, the number of .35 and .20 stamps respectively.\[\{35x+20y=6295,x+y=200\}\]\[\{x=153,y=47\} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

I am having problems putting words into equations Thank You! : D